# SuperPrimitive: Scene Reconstruction at a Primitive Level

### v1

@@ Abstract
We demonstrate the versatility of our new representation by addressing three 3D reconstruction tasks: depth completion, few-view structure from motion, and monocular dense visual odometry.
@@ 1 Introduction
Figure 1: Multi-View Geometry with SuperPrimitives. SuperPrimitives are extracted from an input frame by dividing it into image segments equipped with estimated surface normal directions (bottom-left). Each SuperPrimitive induces a dense reconstruction within the corresponding image segment up to a-priori unknown scale. Different possible reconstructions are shown in light blue. The scales are then jointly optimised together with a relative camera pose to fit multi-view photometric constraints (visualised in green and red). The resulting dense reconstruction of the reference frame is shown in the top.
@@
Our new representation showcases its versatility in three key applications, where geometrical observations from other views have to be incorporated into the prediction:
@@ 3 Method
Our key idea is to estimate local geometry within each segment from a single view, while leaving the relative positioning of the segments to be estimated via multi-view photo-metric consistency optimisation. We refer to these geometrically enhanced minimal image segments as SuperPrimitives, since they are inspired by both superpixels [37] and geometric 3D primitives.
For per-segment local geometry estimation we employ an off-the-shelf surface normal prediction network [1], to estimate infinitesimal geometry for each image pixel. The surface normals within an image segment can be used to estimate its depth via simple integration up to a scale factor. We set these scale factors — depth seeds — to be optimisiable parameters, which are either optimised via multi-view cues or explicit depth seeds for the depth completion experiments.
@@ 3.1 Conventions
Unless stated otherwise, we use lowercase letters for scalar values, bold letters for vectors, and uppercase letters for matrices. We consider images $\mathrm{I}\in\mathbb{R}^{3\times h\times w}$ of height $h$ and width $w$ captured by a camera with a known calibration matrix $\mathrm{K}\in\mathbb{R}^{3\times 3}$. Image pixels are parameterised as $\bm{u}=(u,v)\in\left[0,h-1\right]\times\left[0,w-1\right]$. Given a per pixel depth function $z(\bm{u})\colon\mathrm{I}\to\mathbb{R}^{+}$, we define image’s unprojection onto the 3D space to be $\pi^{-1}(\bm{u})=z(\bm{u})\mathrm{K}^{-1}\bm{\dot{u}}$, where dot is a homogenisation operator $\bm{\dot{p}}=(\bm{p},1)$. Conversely, for a set of 3D points $\bm{x}=(x,y,z)\in\mathbb{R}^{3}$ we define the projection $\pi(\bm{x})=\gamma\left(\frac{1}{z}K\bm{x}\right)$, where the $\gamma$ function drops the last coordinate $\gamma((x,y,z)^{T})=(x,y)^{T}$.
@@ 3.2 SuperPrimitives
We split the input image $\mathrm{I}$ into a set of possibly overlapping connected regions (segments), each of which is equipped with its local 3D geometry. More formally, a SuperPrimitive $P=(\Omega,\mathfrak{D})$ is a connected image region $\Omega\subseteq\mathrm{I}$ which is also equipped with an unscaled depth map $\mathfrak{D}\colon\Omega\to[0,\infty)$ for each pixel within the region $\Omega$. Here, unscaled means that each SuperPrimitive’s depth $\mathfrak{D}$ differs from its ground truth depth $D$ by an a priori unknown scalar. In other words, there exists a single scalar $d$ such that:
$D=d\cdot\mathfrak{D}$ for all pixels p $\text{for all pixels p }\in\Omega$ (1)
@@
For brevity we are use the words primitive and SuperPrimitive interchangeably throughout the rest of the paper.
@@
A set of primitives $\mathcal{P}$ itself is not enough to extract dense image geometry due to scale/depth ambiguity. Therefore, we introduce the concept of optimisable depth seeds which anchor the correct depth scale for each primitive.
Given a scalar depth seed $d$ for a super-primitive $P$, one can infer its depth as $D(p)=d\cdot\mathfrak{D}(p)$ for every point $p\in\Omega$. In practice, we employ a log-depth representation to represent depth throughout the whole system. This means that we store optimisable log-depth seeds $\log d$ and unscaled log-depth $\log\mathfrak{D}$, and so depth inference reduces to a simple shift operation $\log D=\log d+\log\mathfrak{D}$.
In contrast with other dense methods, we estimate the dense geometry $\mathcal{G}$ of the image in the form of a point cloud rather than a depth image. This choice is driven by the fact that multiple primitives’ supporting segments $\Omega_{i}$ and $\Omega_{j}$ might overlap, which would lead to different depth estimates within their intersection $\Omega_{i}\cap\Omega_{j}$.
Given a set of depth seeded primitives =$\mathbb{P}=\{(d_{i},P_{i})\}$, the geometry $\mathcal{G}$ of the image $I$ can be estimated as follows:
$\mathcal{G}\coloneqq\bigcup_{i}\,\pi^{-1}(d_{i}\cdot\mathfrak{D}_{i})$ (2)
@@
Note that this approach is premised on the assumption of geometry being continuous within each predicted image segment $\Omega_{i}$, hence its unscaled depth values could be obtained by integrating its surface normals. Although this is not guaranteed a-priori, we observed that a conservative mask selection demonstrates compelling correlation with geometrical continuity. We discuss this in more detail in the following section.
@@
This ambiguity prompted us to introduce the notion of depth seeds, to allow each segment to be adjusted towards its true depth. We implemented batched normal integration, efficiently solving all optimisation problems as a single sparse linear system using conjugate gradient method [18]. This implementation enables the integration of approximately $100$ to $300$ segments within a total time frame of $\sim\!100$ms.
@@ 3.3.1 Two-view SfM on SuperPrimitives
For the sake of clarity, we formulate our method in the simplest case of two frames observing the scene, although it could be trivially extended into a setting with a higher number of views. At the core of our SfM approach lies a per-primtive photometric alignment, which contrasts with widely accepted per-pixel photometric alignment [30, 14] techniques. Informally speaking, our proposition is to treat each image primitive as a “rigid” piece, which is only allowed to be scaled up or down. This degree of freedom is due to the scale ambiguity present for each primitive. Thus, instead of estimating a depth value per pixel, we only estimate a depth-seed per primitive, which is illustrated in Fig. 1. This greatly reduces the dimensionality of the optimisation problem, especially in the case of an unknown relative pose. Note that our method does not require the target image to be primified nor does it not require any pre-established correspondences (e.g. primitive to primitive).
@@
We assume a primified reference image $\mathrm{I}_{\text{ref}}$ with its set of primitives $\mathcal{P}(\mathrm{I})=\{(\Omega_{i},\mathfrak{D}_{i})\}$. Given an unposed target image $\mathrm{I}_{\text{t}}$, our goal is to jointly estimate its relative pose $\mathrm{T}_{tr}$ with respect to the reference image $\mathrm{I}_{\text{ref}}$ as well as the dense geometry $\mathcal{G}_{\text{ref}}=\bigcup_{i}\,\pi^{-1}(d_{i}\cdot\mathfrak{D}_{i})$ of the reference frame.
@@
While unscaled depths $\mathfrak{D}_{i}$ are given by a per-image pre-processor after the primification of the reference image, the set of depth seeds $d_{i}$ are yet to be estimated. In our case, we jointly estimate both depth seeds ${d_{i}}$ and a relative image pose $\mathrm{T}_{tr}$ by solving a photometric consistency optimisation problem. First, we warp each depth-seeded primitive $(d_{i},P_{i})$ from the reference frame $\mathrm{I}_{\text{ref}}$ into the target frame $\mathrm{I}_{\text{t}}$:
$\hat{P_{i}}[\bm{u}]=\pi\left(\mathrm{T}_{tr}\pi^{-1}\left(\bm{u},d_{i}\cdot\mathfrak{D}_{i}\right)\right)$ (5)
@@
Then a per-segment photometric residual for the primitive $P_{i}=(\Omega_{i},\mathfrak{D})$ is defined by averaging all photometric reprojection $\ell^{1}$ errors across every pixel $\bm{u}\in\Omega_{i}$:
$r(P_{i},d_{i},\mathrm{I}_{\text{t}},\mathrm{T}_{tr})=\frac{1}{|\Omega_{i}|}\sum_{\bm{u}\in\Omega_{i}}\norm{\mathrm{I}_{\text{ref}}(\bm{u})-\mathrm{I}_{\text{t}}(\hat{P_{i}}[\bm{u}])}_{1}$ (6)
@@
The resulting photometric cost, aggregated across all depth-seeded primitives $(d_{i},P_{i})\in\mathcal{P}(\mathrm{I})$, is:
$E_{\text{photo}}=\frac{1}{|\mathcal{P}(\mathrm{I})|}\sum_{P_{i}=(\Omega_{i},\mathfrak{D})}r(P_{i},d_{i},\mathrm{I}_{\text{t}},\mathrm{T}_{tr})$ (7)
@@
Note that we abstain from explicit occlusion handling in our photometric alignment. We, however, expect our system could further be refined with explicit primitive correspondence checks.
@@
Finally, to obtain the relative pose and depth seeds, we minimise the photometric cost using the Adam [20] optimiser:
$\mathrm{T}_{tr}^{\text{opt}},\{d_{i}^{\text{opt}}\}=\operatorname{argmin}\limits_{\mathrm{T}_{tr},d_{i}}\,\,E_{\text{photo}}$ (8)
@@
Our approach trivially extends to multiple reference and target views via photometric cost summation, and we can jointly solve all depth seeds and poses using the resulting cost function.
@@ 3.3.2 Monocular Visual Odometry
Initialisation of a VO system typically involves estimating the geometry of the first keyframe and its pose with respect to the subsequent few frames. Initialisation of a monocular incremental SfM or SLAM system is often hard, because one has to solve a chicken-and-egg problem: poses are required to reliably estimate geometry and vice versa. Therefore many VO / SLAM systems employ a wide range of heuristics in order to initialise their system. However, using our new representation no special treatment is required in our case. In order to bootstrap our system, we simply create the two first keyframes and the employ the Structure-from-Motion method proposed in Sec. 3.3.1. We keep the pose of the first keyframe fixed to fix gauge freedom.
@@
We solve the same photometric cost formulated in Eq. 8, but for the pose only. The latest keyframe serves as the reference frame $\mathrm{I}_{\text{ref}}$ whereas the new incoming image is set to be the target frame $\mathrm{I}_{\text{t}}$. The depths seeds of the latest keyframe have already been estimated at the mapping stage.
@@
The mapping stage ensures geometric consistency within the keframe sliding window of size $n$. Concretely, our mapping refines depth seeds $\{d_{j}\}$ and a pose $\mathrm{T}^{i}_{\text{kf}}$ for each keyframe $\mathrm{I}^{i}_{\text{kf}}$ where $0\leq i\leq n$.
For each keyframe, we define a set of target frames against which the photometric cost is being optimised with a connectivity function $\mathcal{M}(t)=\{\mathrm{I}_{s},\mathrm{T}_{s}\}$. In our implementation, $\mathcal{M}(t)$ includes temporally neighbouring keyframes $\mathrm{I}^{t-1}_{\text{kf}}$ and $\mathrm{I}^{t+1}_{\text{kf}}$ (if they exist) used as supporting frames, as well as $4$ additional supplementary views, for which only the pose is being estimated. We employ these views to better constraint the geometry and observed that increasing the number of supplementary views does not induce high computational cost.
@@
$E_{\text{mapping}}=\sum_{t=0}^{n}\sum_{{\mathrm{I}_{s}\in\mathcal{M}(t)}}\sum_{{P_{i}\in\mathcal{P}({\scriptscriptstyle\mathrm{I}^{t}_{kf}})}}r(P_{i},d_{i},\mathrm{I}_{s},\mathrm{T}^{-1}_{s}\mathrm{T}_{kf})$ (9)
@@ 4.1 Sparse Depth Completion
Figure 3: Depth Completion on VOID. We visualise the coloured unprojections of ground truth depth maps (top row) and the geometry estimated by our method (bottom row). Sparse depth input points are visualised as red dots (electronic zoom-in recommended). Qualitatively, we achieve sharper geometry estimates than from a commodity depth sensor.
Our SuperPrimitives representation could be seamlessly applied to depth completion with no pre-training required, thereby solving it in a zero-shot manner. For each primitive $P_{i}\in\mathcal{P}(\mathrm{I})$ we adjust its depth seed to minimise depth discrepancy $\|d_{i}\cdot\mathfrak{D}_{i}-\hat{D}\|$ with given ground truth sparse depth $\hat{D}$ across all valid depth points within the segment $\Omega_{i}$. If a segment lacks valid depth measurements, it is discarded. For image regions not covered by any valid primitives, we generate a dense depth prediction by simply bilinearly interpolating depth values.
Many existing depth completion studies evaluate their methods in artificial scenarios, such as selecting input depth points randomly from a known ground truth. In contrast, our approach is tested using the real-world VOID benchmark [55]. The benchmark provides video sequences captured with a RealSense D435i camera together with sparse metric depth measurements acquired from an external visual-inertial Structure-from-Motion system. This setup exhibits noise and biases that might be present in the sparse depth inputs. These measurements are already provided by the dataset itself and are therefore shared across all the methods being compared.
@@
In Tab. 1 our method quantitatively performs on par with a recent state-of-the-art method, which is enabled by a monocular depth predictor pre-trained on a vast mixture of datasets [36, 35] and fine-tined on the VOID train set. In the zero-shot setting, we outperform VI-Depth [53], which uses a DPT-Hybrid pre-trained backbone, on three out of four metrics. Note that no training was done the for depth completion task for our method. We also provide a qualitative comparison against ground truth depth in Fig. 3. It could be qualitatively seen that our method provides sharper geometry estimates than the ground truth depth maps, which exhibit noise characteristic to depth sensors.
@@
 $*{}^{*}$ (MiDaS) [53] 113.27 193.38 53.86 84.82
 $*{}^{*}$ (DPT-Hybrid) [53] 97.03 167.82 46.62 74.67
Table 1: Depth Completion on VOID. We report four most widely used metrics for depth completion on the VOID benchmark. The methods which did not use the VOID dataset for training are in the top section. Methods operating in a zero-shot scenario are shown Methods trained on VOID are marked with an $*{}^{*}$. Our method demonstrates superior performance on three our of the four metrics within the zero-shot group. It is a second best on two out the four metrics overall.

### v2

@@ Abstract
We demonstrate the versatility of our new representation by addressing three 3D reconstruction tasks: depth completion, few-view structure from motion, and monocular dense visual odometry. Project page: https://makezur.github.io/SuperPrimitive/
@@ 1 Introduction
Figure 1: Multi-View Geometry with SuperPrimitives. SuperPrimitives are extracted from an input frame by dividing it into image segments equipped with estimated surface normal directions (bottom-left). Each SuperPrimitive induces a dense reconstruction within the corresponding image segment up to a priori unknown scale. Different possible reconstructions are shown in light blue. The scales are then jointly optimised together with a relative camera pose to fit multi-view photometric constraints (visualised in green and red). The resulting dense reconstruction of the reference frame is shown in the top.
@@
Our new representation showcases its versatility in three key applications, where single-view ambiguity is resolved via additional measurements or viewpoints:
@@ 3 Method
Our key idea is to estimate local geometry within each segment from a single view, while leaving the relative positioning of the segments to be estimated via multi-view photometric consistency optimisation. We refer to these geometrically enhanced minimal image segments as SuperPrimitives, since they are inspired by both superpixels [39] and geometric 3D primitives.
For per-segment local geometry estimation we employ an off-the-shelf surface normal prediction network [1], to estimate infinitesimal geometry for each image pixel. The surface normals within an image segment can be used to estimate its depth via simple integration up to a scale factor. We set these scale factors — depth scales — to be optimisiable parameters, which are either optimised via multi-view cues or explicit depth measurements for the depth completion experiments.
@@ 3.1 Conventions
Unless stated otherwise, we use lowercase letters for scalar values, bold letters for vectors, and uppercase letters for matrices. We consider images $\mathrm{I}\in\mathbb{R}^{3\times h\times w}$ of height $h$ and width $w$ captured by a camera with a known calibration matrix $\mathrm{K}=\begin{pmatrix}f_{u}&0&c_{u}\\ 0&f_{v}&c_{v}\\ 0&0&1\end{pmatrix}\in\mathbb{R}^{3\times 3}$. Image pixels are parameterised as $\bm{u}=(u,v)\in\left[0,h-1\right]\times\left[0,w-1\right]$. Given a per pixel depth function $z(\bm{u})\colon\mathrm{I}\to\mathbb{R}^{+}$, we define image’s unprojection onto the 3D space to be $\pi^{-1}(\bm{u})=z(\bm{u})\mathrm{K}^{-1}\bm{\dot{u}}$, where dot is a homogenisation operator $\bm{\dot{p}}=(\bm{p},1)$. Conversely, for a set of 3D points $\bm{x}=(x,y,z)\in\mathbb{R}^{3}$ we define the projection $\pi(\bm{x})=\gamma\left(\frac{1}{z}K\bm{x}\right)$, where the $\gamma$ function drops the last coordinate $\gamma((x,y,z)^{T})=(x,y)^{T}$.
@@ 3.2 SuperPrimitives
We split the input image $\mathrm{I}$ into a set of possibly overlapping connected regions (segments), each of which is equipped with its local 3D geometry. More formally, a SuperPrimitive $P=(\Omega,\mathfrak{D})$ is a connected image region $\Omega\subseteq\mathrm{I}$ which is also equipped with an unscaled depth map $\mathfrak{D}\colon\Omega\to[0,\infty)$ for each pixel within the region $\Omega$. Here, unscaled means that each SuperPrimitive’s depth $\mathfrak{D}$ differs from its ground truth depth $D$ by an a priori unknown scalar. In other words, there exists a single scalar $s$ such that:
$D=s\cdot\mathfrak{D}$ for all pixels p $\text{for all pixels p }\in\Omega$ (1)
@@
For brevity we use the words primitive and SuperPrimitive interchangeably throughout the rest of the paper.
@@
A set of primitives $\mathcal{P}$ itself is not enough to extract dense image geometry due to scale/depth ambiguity. Therefore, we introduce the concept of optimisable depth scales which anchor the correct depth scale for each primitive.
Given a scalar depth scale $s$ for a super-primitive $P$, one can infer its depth as $D(p)=s\cdot\mathfrak{D}(p)$ for every point $p\in\Omega$. In practice, we employ a log-depth representation to represent depth throughout the whole system. This means that we store optimisable log-depth scales $\log s$ and unscaled log-depth $\log\mathfrak{D}$, and so depth inference reduces to a simple shift operation $\log D=\log s+\log\mathfrak{D}$.
In contrast to other dense methods, we estimate the dense geometry $\mathcal{G}$ of the image in the form of a point cloud rather than a depth image. This choice is driven by the fact that multiple primitives’ supporting segments $\Omega_{i}$ and $\Omega_{j}$ might overlap, which would lead to different depth estimates within their intersection $\Omega_{i}\cap\Omega_{j}$.
Given a set of depth scaled primitives =$\mathbb{P}=\{(s_{i},P_{i})\}$, the geometry $\mathcal{G}$ of the image $I$ can be estimated as follows:
$\mathcal{G}\coloneqq\bigcup_{i}\,\pi^{-1}(s_{i}\cdot\mathfrak{D}_{i})$ (2)
@@
Note that this approach is premised on the assumption of geometry being continuous within each predicted image segment $\Omega_{i}$, hence its unscaled depth values could be obtained by integrating its surface normals. Although this is not guaranteed a priori, we observed that a more fine-grained mask selection yields a compelling correlation with geometrical continuity. We discuss this in more detail in the following section.
@@
This ambiguity prompted us to introduce the notion of depth scales, to allow each segment to be adjusted towards its true depth. We implemented batched normal integration, efficiently solving all optimisation problems as a single sparse linear system using conjugate gradient method [18]. This implementation enables the integration of approximately $100$ to $300$ segments within a total time frame of $\sim\!100$ms.
@@ 3.3.1 Two-view SfM on SuperPrimitives
For the sake of clarity, we formulate our method in the simplest case of two frames observing the scene, although it could be trivially extended into a setting with a higher number of views. At the core of our Structure-from-Motion (SfM) approach lies a per-primtive photometric alignment, which contrasts with widely accepted per-pixel photometric alignment [30, 14] techniques. Informally speaking, our proposition is to treat each image primitive as a “rigid” piece, which is only allowed to be scaled up or down. This degree of freedom is due to the scale ambiguity present for each primitive. Thus, instead of estimating a depth value per pixel, we only estimate a depth-scale per primitive, which is illustrated in Fig. 1. This greatly reduces the dimensionality of the optimisation problem, especially in the case of an unknown relative pose. Note that our method does not require the target image to be primified nor does it not require any pre-established correspondences (e.g. primitive to primitive).
@@
We assume a primified reference image $\mathrm{I}_{\text{ref}}$ with its set of primitives $\mathcal{P}(\mathrm{I})=\{(\Omega_{i},\mathfrak{D}_{i})\}$. Given an unposed target image $\mathrm{I}_{\text{t}}$, our goal is to jointly estimate its relative pose $\mathrm{T}_{tr}$ with respect to the reference image $\mathrm{I}_{\text{ref}}$ as well as the dense geometry $\mathcal{G}_{\text{ref}}=\bigcup_{i}\,\pi^{-1}(s_{i}\cdot\mathfrak{D}_{i})$ of the reference frame.
@@
While unscaled depths $\mathfrak{D}_{i}$ are given by a per-image pre-processor after the primification of the reference image, the set of depth scales $s_{i}$ are yet to be estimated. In our case, we jointly estimate both depth scales ${s_{i}}$ and a relative image pose $\mathrm{T}_{tr}$ by solving a photometric consistency optimisation problem. First, we warp each depth-scaled primitive $(s_{i},P_{i})$ from the reference frame $\mathrm{I}_{\text{ref}}$ into the target frame $\mathrm{I}_{\text{t}}$:
$\hat{P_{i}}[\bm{u}]=\pi\left(\mathrm{T}_{tr}\pi^{-1}\left(\bm{u},s_{i}\cdot\mathfrak{D}_{i}\right)\right)$ (5)
@@
Then a per-segment photometric residual for the primitive $P_{i}=(\Omega_{i},\mathfrak{D}_{i})$ is defined by averaging all photometric reprojection $\ell^{1}$ errors across every pixel $\bm{u}\in\Omega_{i}$:
$r(P_{i},s_{i},\mathrm{I}_{\text{t}},\mathrm{T}_{tr})=\frac{1}{|\Omega_{i}|}\sum_{\bm{u}\in\Omega_{i}}\norm{\mathrm{I}_{\text{ref}}(\bm{u})-\mathrm{I}_{\text{t}}(\hat{P_{i}}[\bm{u}])}_{1}$ (6)
@@
The resulting photometric cost, aggregated across all depth-scaled primitives $(s_{i},P_{i})\in\mathcal{P}(\mathrm{I})$, is:
$E_{\text{photo}}=\frac{1}{|\mathcal{P}(\mathrm{I})|}\sum_{P_{i}=(\Omega_{i},\mathfrak{D})}r(P_{i},s_{i},\mathrm{I}_{\text{t}},\mathrm{T}_{tr})$ (7)
@@
Note that we abstain from explicit occlusion handling in our photometric alignment, as our per-segment alignment is by design robust to pixel-level occlusion. We, however, expect our system could further be refined with explicit primitive correspondence checks.
@@
Finally, to obtain the relative pose and depth scales, we minimise the photometric cost using the Adam [20] optimiser:
$\{s_{i}^{\text{opt}},\mathrm{T}_{tr}^{\text{opt}}\}=\operatorname{argmin}\limits_{s_{i},\mathrm{T}_{tr}}\,\,E_{\text{photo}}$ (8)
@@
Our approach trivially extends to multiple reference and target views via photometric cost summation, and we can jointly solve all depth scales and poses using the resulting cost function.
@@ 3.3.2 Monocular Visual Odometry
Initialisation of a VO system typically involves estimating the geometry of the first keyframe and its pose with respect to the subsequent few frames. Initialisation of a monocular incremental SfM or SLAM system is often hard, because one has to solve a chicken-and-egg problem: poses are required to reliably estimate geometry and vice versa. Therefore many VO / SLAM systems employ a wide range of heuristics in order to initialise their system. However, using our new representation no special treatment is required in this case. In order to bootstrap our system, we simply create the two first keyframes and then employ the Structure-from-Motion method proposed in Sec. 3.3.1. We keep the pose of the first keyframe fixed to fix gauge freedom.
@@
We solve the same photometric cost formulated in Eq. 8, but for the pose only. The latest keyframe serves as the reference frame $\mathrm{I}_{\text{ref}}$ whereas the new incoming image is set to be the target frame $\mathrm{I}_{\text{t}}$. The depth scales of the latest keyframe have already been estimated at the mapping stage.
@@
The mapping stage ensures geometric consistency within the keyframe sliding window of size $n$. Concretely, our mapping refines depth scales $\{s_{j}\}$ and a pose $\mathrm{T}^{i}_{\text{kf}}$ for each keyframe $\mathrm{I}^{i}_{\text{kf}}$ where $0\leq i\leq n$.
For each keyframe, we define a set of target frames against which the photometric cost is being optimised with a connectivity function $\mathcal{M}(t)=\{\mathrm{I}_{s},\mathrm{T}_{s}\}$. In our implementation, $\mathcal{M}(t)$ includes temporally neighbouring keyframes $\mathrm{I}^{t-1}_{\text{kf}}$ and $\mathrm{I}^{t+1}_{\text{kf}}$ (if they exist) used as supporting frames, as well as $4$ additional supplementary views, for which only the pose is being estimated. We employ these views to better constrain the geometry and observed that increasing the number of supplementary views does not induce high computational cost.
@@
$E_{\text{mapping}}=\sum_{t=0}^{n}\sum_{{\mathrm{I}_{s}\in\mathcal{M}(t)}}\sum_{{P_{i}\in\mathcal{P}({\scriptscriptstyle\mathrm{I}^{t}_{kf}})}}r(P_{i},s_{i},\mathrm{I}_{s},\mathrm{T}^{-1}_{s}\mathrm{T}_{kf})$ (9)
@@ 4.1 Sparse Depth Completion
Figure 3: Depth Completion on VOID. We visualise the coloured unprojections of ground truth depth maps provided by a sensor (top row) and the geometry estimated by our method (bottom row). Sparse depth input points are visualised as red dots (electronic zoom-in recommended). Qualitatively, we achieve sharper geometry estimates than from a commodity depth sensor.
Our SuperPrimitives representation could be seamlessly applied to depth completion with no pre-training required, thereby solving it in a zero-shot manner. For each primitive $P_{i}\in\mathcal{P}(\mathrm{I})$ we adjust its depth scale to minimise depth discrepancy $\|s_{i}\cdot\mathfrak{D}_{i}-\hat{D}\|$ with given ground truth sparse depth $\hat{D}$ across all valid depth points within the segment $\Omega_{i}$. If a segment lacks valid depth measurements, it is discarded. For image regions not covered by any valid primitives, we generate a dense depth prediction by simply bilinearly interpolating depth values.
Many existing depth completion studies evaluate their methods in artificial scenarios, such as selecting input depth points randomly from a known ground truth. In contrast, our approach is tested using the real-world VOID benchmark [57]. The benchmark provides video sequences captured with a RealSense D435i camera together with sparse metric depth measurements acquired from an external visual-inertial SfM system. This setup exhibits noise and biases that might be present in the sparse depth inputs. These measurements are already provided by the dataset itself and are therefore shared across all the methods being compared.
@@
In Tab. 1 our method quantitatively performs on par with a recent state-of-the-art method, which is enabled by a monocular depth predictor pre-trained on a vast mixture of datasets [38, 37] and fine-tined on the VOID train set. In the zero-shot setting, we outperform VI-Depth [55], which uses a DPT-Hybrid pre-trained backbone, on three out of four metrics. Note that no training was done the for depth completion task for our method. Compared to the ground truth depth maps obtained via a noisy sensor, our predictions show (Fig. 3) sharper object boundaries and are better at preserving straight lines and perpendicular structures.
@@
VI-Depth (MiDaS) [55] 113.27 193.38 53.86 84.82
VI-Depth (DPT-Hybrid) [55] 97.03 167.82 46.62 74.67
Table 1: Depth Completion on VOID. We report four most widely used metrics for depth completion on the VOID benchmark. The methods which did not use the VOID dataset for training are in the top section. Methods trained on VOID are marked with an asterisk. Our method demonstrates superior performance on three out of the four metrics within the zero-shot group. It is second-best on two out of the four metrics overall.